OpenStudy (anonymous):

remainder of 3^1990 : 41 is ...

ganeshie8 (ganeshie8):

use below : 3^4 leaves a remainder -1

OpenStudy (anonymous):

1990 is disable divided by 41, @ganeshi

ganeshie8 (ganeshie8):

\(3^{1990} = 3^{497\times 4 + 2} = 3^2 3^{497\times 4} = 9 (3^4)^{497}\)

ganeshie8 (ganeshie8):

since \(3^4\) leaves a remainder \(-1\), \((3^4)^{497}\) leaves a remainder \((-1)^{497}\)

OpenStudy (anonymous):

honestly, i m not good in modulo. i dont know which properties i should use

OpenStudy (anonymous):

can you show to me step by step to solve this, @ganeshie8

ganeshie8 (ganeshie8):

we need to use only two properties : 1) if \(a \) leaves a remainder \(r\), then \(a^n\) leaves a remainder \(r^n\) 1) if \(a \) leaves a remainder \(r\), then \(a\times c\) leaves a remainder \(a\times r\)

ganeshie8 (ganeshie8):

you familiar wid congruence notation right ?

OpenStudy (anonymous):

yeah, it like 3^1990 mod(41) = ...

ganeshie8 (ganeshie8):

corrected typo.. we need to use only two properties : 1) if \(a\) leaves a remainder r, then an leaves a remainder \(r^n\) 2) if \(a\) leaves a remainder r, then a×c leaves a remainder r×c

ganeshie8 (ganeshie8):

ahh then its easy

ganeshie8 (ganeshie8):

\(3^{1990} \equiv 3^{4\times 497 + 2} \equiv 3^2 (3^4)^{497}\mod 41 \)

ganeshie8 (ganeshie8):

fine so far ? we hvaent used any mod properties yet... just broke down the exponent...

OpenStudy (anonymous):

ok, so far i got 9(81)^497 mod(41) what's next ? btw, can you type the equation more big before ?, thanks :)

ganeshie8 (ganeshie8):

\(\large 3^{1990} \equiv 3^{4\times 497 + 2} \equiv 3^2 (3^4)^{497}\mod 41 \)

ganeshie8 (ganeshie8):

this looks better ?

OpenStudy (anonymous):

yeah, that's better than before :)

ganeshie8 (ganeshie8):

next, whats the remainder when 3^4 is divided by 41 ?

OpenStudy (anonymous):

40

OpenStudy (anonymous):

81/41 = 1, remainder 40 right ?

ganeshie8 (ganeshie8):

yes, 81/41 : 40 is the overflow also, -1 is the underflow

ganeshie8 (ganeshie8):

saying "40" is the remainder is same as saying "-1" is the remainder

ganeshie8 (ganeshie8):

40 - 41 = -1

OpenStudy (anonymous):

ah, yes i know because 81-(-1) is multiply of 41, right ?

ganeshie8 (ganeshie8):

yes, thats more elegant way of putting it :)

ganeshie8 (ganeshie8):

go back to the mod

ganeshie8 (ganeshie8):

\(\large 3^{1990} \equiv 3^{4\times 497 + 2} \equiv 3^2 (3^4)^{497} \equiv 9(-1)^{497} \mod 41\)

ganeshie8 (ganeshie8):

see if that looks still fine

OpenStudy (anonymous):

(-1)^497 = -1 so, 9(-1) mod 41 = -9 mod 41 = ... what's next ?

ganeshie8 (ganeshie8):

we're done

ganeshie8 (ganeshie8):

-9 is the remainder

ganeshie8 (ganeshie8):

-9 is same as -9+41

OpenStudy (anonymous):

huh ? -9, it should be positive ?

ganeshie8 (ganeshie8):

just add 41 if u want positive answer

OpenStudy (anonymous):

ohhhhhhhhhhhh.... im very stupid with this, that's make sense now. before i used binomial newtown's theorem, twhere to longg. but i got same answer. it is 32

OpenStudy (anonymous):

thank you very much, @ganeshie8 . btw, do you know the sites where i get more question to training my self ?

ganeshie8 (ganeshie8):

oh interesting, i thought of binomial in the start... but couldn't figure out how the simplification is possible using binomial...

ganeshie8 (ganeshie8):

if its not lengthy, may i knw how did u proceed wid the binomial thm ?

OpenStudy (anonymous):

same like the 1st. 3^(1990) : 41 = (3)^(4*497 + 2) : 41 = 9 * (3^4)^497 : 41 = 9 * (81)^497 : 41 = 9 * (82-1)^497 : 41 see, that (82-1)^497 = 497C0 (82)^497 - 497C1 (82)^496 + 497C2 (82)^495 + ... + 497C495(82)^2(-1) + 497C496 (82)(-1)^498 + 497C497 (-1)^497 i assumed that 497C0 (82)^497 - 497C1 (82)^496 + 497C2 (82)^495 + ... - 497C495(82)^2(-1), is P. where obvious this sum able divided by 41. so, 9 * (82-1)^497 : 41 = 9 * [P + 497C496 (82)(-1)^498 + 497C497 (-1)^497] : 41 = 9 * [P + 497 * 82 - 1] : 41 = 9 * [P + 40,753] : 41 = (9P + 9*40,753) : 41 = 9P/41 + (9*40,753 : 41) (9*40,753 : 41), from here i got the remainder 32 :)

ganeshie8 (ganeshie8):

Excellent !! basically u have simplified it to form : 9 * (82-1)^497 = 9( 82*something - 1) since 82 is divisible by 41, remainder is -9

OpenStudy (anonymous):

but looks, my way is stupid, very-very stupid. too longggggggg :V

ganeshie8 (ganeshie8):

noo, its not long. try below when you're free : find the remainder when 5^99 is divided by 13

ganeshie8 (ganeshie8):

we can do this both ways : mods, and binomial

OpenStudy (anonymous):

well, i'll try this :)

OpenStudy (anonymous):

5^99 mod(13) = 5^(2*49 + 1) mod(13) = 5 * 5^(2*49) mod (13) = 5 * 25^49 mod (13) = 5 * (26-1)^49 mod (13) wait.. what should i do now

ganeshie8 (ganeshie8):

\(25 \equiv -1 \mod 13 \)

ganeshie8 (ganeshie8):

5^99 mod(13) = 5^(2*49 + 1) mod(13) = 5 * 5^(2*49) mod (13) = 5 * 25^49 mod (13) = 5 * (-1)^49 mod (13)

ganeshie8 (ganeshie8):

binomial thm is also easy here..

OpenStudy (anonymous):

5 * (26 * something - 1) since 26 is divisible by 13, remainder is -5 haha.. i copied your statement :)

ganeshie8 (ganeshie8):

yup :) both methods give answer in one step !!

ganeshie8 (ganeshie8):

here more problems : http://www.mathdb.org/notes_download/elementary/number/ne_N2.pdf

OpenStudy (anonymous):

so, then just add -5 with 13 ?

ganeshie8 (ganeshie8):

yes !

ganeshie8 (ganeshie8):

8 is the remainder

ganeshie8 (ganeshie8):

try Example 1.1 problems in above link when ur free... have fun :)

OpenStudy (anonymous):

waaw, math is beatiful. i dont know before, if i get negative then just add with divisor

ganeshie8 (ganeshie8):

number theory is beautiful and easy if we use intuition enough :) yes, also saying "-5" is remainder is perfectly fine as well. more on this here : http://openstudy.com/users/ganeshie8#/updates/533ef9e1e4b05a4f7d53c3ee

OpenStudy (anonymous):

thank you so much, for your helps and for the links. i;ll try those later ... :)

ganeshie8 (ganeshie8):

np :)

OpenStudy (anonymous):

just wanna quick confirmation are you trying to find the remainder of \[3^{1990} (\mod 41)\]? because i got a different remainder

OpenStudy (anonymous):

yeah, i got 32

OpenStudy (anonymous):

oh sorry nvm i looked i the wrong posts lol yea i got 32 as well i wanna gonna suggest using Fermat's Little Theorem to do this problem

ganeshie8 (ganeshie8):

^Fermat is the fastest way

OpenStudy (anonymous):

To confirm your answer visit http://www.wolframalpha.com/input/?i=remainder+%28+3^1990%2C+41%29